OpenStudy (anonymous):

determine a system of equations in three variables that models the problem, then solve: the sum of three numbers is nine. two times the first number plus the second number less the third number is 30. three times the first number subtract four times the second number plus two times the third is -49.

OpenStudy (anonymous):

Start with giving all the numbers variables a, b and c.

OpenStudy (anonymous):

Then write all the equations and state that they are simultaneous.

OpenStudy (anonymous):

As Zuige explained a + b + c = 9 2a + b - c = 30 3a -4b + 2c = -49

OpenStudy (anonymous):

a + b + c = 9 2a + b - c = 30 3a - 4b + 2c = -49

OpenStudy (anonymous):

Then just solve simultaneous equations.

OpenStudy (anonymous):

thank you!